Elimination -2X-3y=5 -5X+3y=-40
Answer by Boreal(15235)

-2X-3y=5
-5X+3y=-40
add them, and the ys disappear
-7x=-35
divide by -7 both sides
x=5
substitute into first equation to get y
(-2)(5)-3y=5
-10-3y=5
-3y=15
y=-5
substitute into second equation to check
-5(5)+3(-5)=-40 checks
(5,-5)